Session: Feb 2005 (session 1)
Wavelength: 21 cm


To give feedback on this experiment, click here

Experiment Feedback

Urumqi (yang wen jun at Fri Mar 4 17:44:31 2005) Success :
The T-sys of RCP is too high,the reason is not clear.the experiment was stopped at UT 17:17:00 because of the following source is outside dish limits.

Medicina (Monari at Fri Mar 4 23:33:02 2005) Minor failures : FS
Some problem during the recording for communication between FS and MV

Onsala (Michael Lindqvist at Sat Mar 5 08:34:29 2005) Success :
No known problems. We have fixed the cooling problem, using cooled HEMTs again.

Torun (K. Borkowski at Mon Mar 7 08:47:17 2005) Minor failures : RFI
BBC3 unlocked all the time (problem not related to the BBC unit itself, but electronics prior to it). Intermittent strong RFI.

Jodrell_Bank (Alastair Gunn at Mon Mar 7 12:19:30 2005) Minor failures : WIND
The Lovell Telescope was used. Antenna was parked for high winds between 11:24 and 13:12 UT. For first 1.5 hours, phase CAL was intermittently on and off, problem solved during parked time. No other known problems.

Noto (P. Cassaro at Mon Mar 7 15:41:05 2005) Minor failures :
The procedure checkmk5a has not been recognized for the whole experiment. It is not present in the experiment procedured nor in the station procedures.

Effelsberg (A. Kraus at Fri Mar 11 12:40:38 2005) Minor failures : WIND
Snowstop: UT1000-1052.